# The discrete periodic Pitman transform: invariances, braid relations, and Burke properties

### v2

@@ Abstract.
We develop the theory of the discrete periodic Pitman transform, first introduced by Corwin, Gu, and the fifth author. We prove that, for polymers in a periodic environment, single-path and multi-path partition functions are preserved under the action of this transform on the weights in the polymer model. As a corollary, we prove that the discrete periodic Pitman transform satisfies the same braid relations that are satisfied for the full-line Pitman transform shown by Biane, Bougerol, and O’Connell. Combined with a new inhomogeneous Burke property for the periodic Pitman transform, we prove a multi-path invariance result for the periodic inverse-gamma polymer under permutations of the column parameters. In the limit to the full-line case, we obtain a multi-path extension of a recent invariance result of Bates, Emrah, Martin, Seppäläinen, and the fifth author, in both positive and zero-temperature.
@@ 1. Introduction
The study of the Pitman transform traces back to Pitman’s celebrated 2MX2M-X theorem [Pit75]. Since then, there have been numerous studies around this transform and its many variants. In the semi-discrete case, the positive-temperature lifting of the full-line Pitman transform was introduced by [MY01, O’C03] in the context of Brownian queuing process and later by [O’C12] in the setting of semi-discrete directed polymers. In the fully discrete case, the discrete geometric Pitman transform was connected to the geometric lifting of the classical Robinson-Schensted-Knuth (RSK) correspondence [Kir01] and studied by [NY04, COSZ14, OSZ14, Cor21] in the context of discrete directed polymer models.
@@
Despite arising in different contexts, these Pitman transforms share a rich set of both algebraic and probabilistic properties that have enabled significant progress in the study of stochastic integrable systems, including directed percolation, random polymers, the KPZ equation, and so on. For example, the Burke property–the analogous product measure preservation property of Burke’s theorem for $M/M/1$ queues [Bur56]–has been proved in several contexts (see [HB76, OY01, OY02, KOR02, Mar09, DMO05, AD95] as well as [EJS24, BEM+25a] in an inhomogeneous setting) and has been used to determine the fluctuation exponents in several last-passage and polymer settings [CG06, BCS06, SV10, Sep12, EJS23, LNS23]. Moreover, the joint invariance of last-passage values/polymer partition functions under the zero-temperature/positive-temperature Pitman transforms provides key tools to the construction of central limiting objects in the KPZ universality class [DOV22] and for proving convergence of many different models to these limiting objects [DV21, Wu23, ACH24, Zha25].
@@
The purpose of this paper is to demonstrate the rich algebraic and probabilistic structure of a periodic version of the Pitman transform, recently introduced in [CGS24]. Our first main result, Theorem 1.4, gives an invariance result for directed polymers in a periodic environment under a periodic Pitman transformation of the column weights. Previous results of this flavor in a non-periodic setting have appeared in several contexts and are discussed after the statement of the theorem. The key machinery is the encoding of geometric RSK in terms of matrix relations from [NY04]. Corwin [Cor21] previously used these tools to prove the analogous invariance for full-space polymers. The next main result, Theorem 1.6, uses this polymer invariance result to prove braid relations for the periodic Pitman transform, analogous to the relations enjoyed by the full-line Pitman transform [BBO05]. More specifically, we consider a collection of operators on the space $(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ of bi-infinite sequences of $N$-periodic vectors and operators $\mathcal{P}_{k}$, which replace the $k$th and k+1k+1st vector with their Pitman transform. The braid relations, along with the fact that each $\mathcal{P}_{k}$ is an involution (proved in [CGS24], see Lemma A.2) shows that this collection of operators generates a copy of the infinite symmetric group. We then utilize the invariance of partition functions and a new inhomogeneous Burke property (Proposition 4.1) to prove a distributional invariance of the inverse-gamma polymer partition function under the permutation of the row and column parameters in both the periodic and full-line cases (Theorems 1.7 and 1.8). In the full-line case, this gives a multi-path and positive temperature extension of a recent result of Bates, Emrah, Martin, Seppäläinen, and the fifth author [BEM+25b]. This result bears similarity to an invariance found previously by Dauvergne [Dau22]. See the discussion before and after Theorem 1.8 for more on the connection to other works. Each of our main results has an analogue in zero-temperature. These results are stated in Section 1.6.
@@ 1.1. The positive-temperature discrete periodic Pitman transform
For $N\in\mathbb{N}$, let $\mathbb{R}^{\mathbb{Z}_{N}}$ denote the set of $N$-tuples, together with periodic indexing. That is, for $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$, $\mathbf{X}=(X_{1},X_{2},\ldots,X_{N})$, where the indices represent elements of $\mathbb{Z}_{N}$ with addition modulo $N$ (see Figure 1). For a vector $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$ and indices $i,j\in\mathbb{Z}_{N}$, we set
@@
where the sums are taken in cyclic order. We follow the convention that $\mathbf{X}_{(i,i]}=0$.
@@
$\mathcal{P}(\mathbf{W}_{1},\mathbf{W}_{2})$ $=(T(\mathbf{W}_{2},\mathbf{W}_{1}),D(\mathbf{W}_{2},\mathbf{W}_{1})),\quad\text{where}$ (1.2)
$T(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{1,i-1}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i,j]}}}\Biggr{)},\quad\text{and}$
$D(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i+1}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}\Biggr{)},\quad\text{where}\quad Y_{\ell}=X_{2,\ell+1}-X_{1,\ell}.$
This map is a shifted version of the map studied in [CGS24]; see Appendix A for the precise connection. Note that we reverse the order of the inputs $\mathbf{W}_{1},\mathbf{W}_{2}$ in the application of $T$ and $D$. Also, note that the maps $\mathcal{P},T,D$ depend on the length $N$ of the periodicity; we will subsume dependence on this value in the notation and fix $N$, unless otherwise noted.
@@ 1.2. Polymers on $\mathbb{Z}^{2}$ and periodic environments
$\pi:=\left\{(a_{i},b_{i})_{i\in\llbracket 1,L\rrbracket}\mid(a_{i+1},b_{i+1})-(a_{i},b_{i})\in\{(1,0),(0,1)\}\ \forall 1\leq i<L\right\}$ (1.3)
where $(a_{i},b_{i})\in\mathbb{Z}^{2}\ \forall i$ and the notation $\llbracket a,b\rrbracket$ denotes the set of $k\in\mathbb{Z}$ satisfying $a\leq k\leq b$. We also define the set of all up-right paths to be $\Pi_{\mathbb{Z}^{2}}$.
@@
to be the set of paths starting at $(i,j)$ and ending at $(n,m)$, noting $\Pi_{(i,j),(n,m)}\neq\varnothing\iff m\geq j$ and $n\geq i$. In this case, L=m+njiL=m+n-j-i for any $\pi$ in $\Pi_{(i,j),(n,m)}$. We say that two paths $\pi_{1},\pi_{2}\in\Pi_{\mathbb{Z}^{2}}$ are non-intersecting if $\pi_{1}\cap\pi_{2}=\varnothing$.
Also of interest to us are sets of non-intersecting paths on our graph. We define a multi-path $(\pi_{1},\ldots,\pi_{k})$ to be a non-intersecting $k$-tuple of paths. See Figure 2 for a visual example.
@@
$\Pi_{U,V}:=\{\pi=(\pi_{1},\dots,\pi_{k})\mid\pi_{i}\in\Pi_{(a_{i},b_{i}),(c_{i},d_{i})}\ \forall i\in\llbracket 1,k\rrbracket,\ \pi_{i}\cap\pi_{j}=\varnothing\ \forall i\neq j\in\llbracket 1,k\rrbracket\}$ (1.5)
We will often require that there is a unique pairing of initial points in $U$ and end points in $V$ such that a non-intersecting multi-path exists. In other words, we require that if $U$ and $V$ each have $k$ elements, then there does not exist a nontrivial permutation $\sigma$ in the $k$th-order symmetric group $S_{k}$ such that $\Pi_{U,V_{\sigma}}\neq\varnothing$ where $V_{\sigma}:=\bigl{(}(c_{\sigma(i)},d_{\sigma(i)})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$. We denote the pairs of $k$-tuples of ordered pairs satisfying this property as
$\Psi:=\bigl{\{}(U,V)\mid U,V\subseteq\mathbb{Z}^{2},\ |U|=|V|,\ \Pi_{U,V}\neq\varnothing,\ \text{ and }\Pi_{U,V_{\sigma}}=\varnothing\ \forall\sigma\in S_{|U|}\setminus\{\text{Id}\}\bigr{\}}$ (1.6)
@@
For real-valued weights $\mathbf{W}=\bigl{(}\mathbf{W}_{i}\in\mathbb{R}^{\mathbb{Z}}\bigr{)}_{i\in\mathbb{Z}}=\bigl{(}W_{(i,j)}\in\mathbb{R}\mid i,j\in\mathbb{Z}\bigr{)}$ corresponding to the vertices of the directed lattice $\mathbb{Z}^{2}$, we define the polymer partition function with inverse temperature $\beta>0$ between the initial point $(i,j)$ and endpoint $(n,m)$ in $\mathbb{Z}^{2}$ as
$Z^{\beta,\mathbf{W}}(n,m\mid i,j):=\sum_{\pi\in\Pi_{(i,j),(n,m)}}\prod_{(r,\ell)\in\pi}e^{\beta W_{(r,\ell)}}.$ (1.7)
@@
$Z^{\beta,\mathbf{W}}(V\mid U):=\sum_{\pi\in\Pi_{U,V}}\prod_{(r,\ell)\in\pi}e^{\beta W_{(r,\ell)}}.$ (1.8)

### v3

@@ Abstract.
We develop the theory of the discrete periodic Pitman transform, first introduced by Corwin, Gu, and the fifth author. We prove that the discrete periodic Pitman transform satisfies the same braid relations that are satisfied for the full-line Pitman transform shown by Biane, Bougerol, and O’Connell. This defines a group action of the infinite symmetric group on sequences of vectors in $\mathbb{R}^{\mathbb{Z}_{N}}$. We prove that, for polymers in a periodic environment, single-path and multi-path partition functions are preserved under the action of this transform on the weights in the polymer model. Combined with a new inhomogeneous Burke property for the periodic Pitman transform, we prove a multi-path invariance result for the periodic inverse-gamma polymer under permutations of the column parameters. In the limit to the full-line case, we obtain a multi-path extension of a recent invariance result of Bates, Emrah, Martin, Seppäläinen, and the fifth author, in both positive and zero-temperature.
@@ 1. Introduction
The study of the Pitman transform traces back to Pitman’s celebrated 2MX2M-X theorem [PIT75]. Since then, there have been numerous studies around this transform and its many variants. In the semi-discrete case, the positive-temperature lifting of the full-line Pitman transform was introduced by [MY01, O’C03] in the context of Brownian queuing process and later by [O’C12] in the setting of semi-discrete directed polymers. In the fully discrete case, the discrete geometric Pitman transform was connected to the geometric lifting of the classical Robinson-Schensted-Knuth (RSK) correspondence [KIR01] and was studied in [NY04, COS+14, OSZ14, COR02] in the context of discrete directed polymer models.
@@
Despite arising in different contexts, these Pitman transforms share a rich set of algebraic and probabilistic properties that have enabled significant progress in the study of stochastic integrable systems, including directed percolation, random polymers, and the KPZ equation. For example, the Burke property–the analogous product measure preservation property of Burke’s theorem for $M/M/1$ queues [BUR56]–has been proved in several contexts (see [HB76, OY01, OY02, KOR02, MAR09, DMO05, AD95] as well as [EJS25, BEM+25] in an inhomogeneous setting) and has been used to compute the fluctuation exponents for several last-passage percolation and polymer models [CG06, BCS06, SV10, SEP12, EJS23, LNS23]. Moreover, the joint invariance of last-passage values and polymer partition functions under the zero-temperature and positive-temperature Pitman transforms provides key tools for constructing central limiting objects in the KPZ universality class [DOV22b] and for proving convergence of many different models to these limiting objects [DV21, WU23, ACH24, ZHA25].
The purpose of this paper is to demonstrate the rich algebraic and probabilistic structure of a periodic version of the Pitman transform, recently introduced in [CGS26]. Our first main result, Theorem 1.1 proves that the periodic Pitman transform is an involution and proves braid relations for these transforms, analogous to the relations enjoyed by the full-line Pitman transform [BBO05] (see also [DNV22a, DV24]). More specifically, we consider a collection of operators on the space $(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ of bi-infinite sequences of $N$-periodic vectors and operators $\mathcal{P}_{k}$, which replace the $k$th and k+1k+1st vector with their Pitman transform. These relations show that this collection of operators generates a copy of the infinite symmetric group, which defines a group action on the space of infinite sequences of vectors in $\mathbb{R}^{\mathbb{Z}_{N}}$.
Our next main result, Theorem 1.4, shows that single-path and multi-path partition functions for directed polymers in a periodic environment are preserved under this group action on the columns of weights for the polymer. Previous results of this flavor in a non-periodic setting have appeared in several contexts and are discussed after the statement of the theorem. The key machinery is the encoding of geometric RSK in terms of matrix relations from [NY04]. Corwin [COR02] previously used these tools to prove the analogous invariance for full-space polymers.
@@
We then utilize the invariance of partition functions and a new inhomogeneous Burke property (Proposition 4.1) to prove a distributional invariance of the inverse-gamma polymer partition function under the permutation of the row and column parameters in both the periodic and full-space cases (Theorems 1.5 and 1.6). In the full-line case, this gives a multi-path and positive-temperature extension of a recent result of Bates, Emrah, Martin, Seppäläinen, and the fifth author [BEM+25]. This result bears similarity to an invariance found previously by Dauvergne [DAU22]. See the discussion before and after Theorem 1.6 for more on the connection to other works. Each of our main results has an analogue in zero-temperature. These are stated in Section 1.5.
@@ 1.1. The positive-temperature discrete periodic Pitman transform
For $N\in\mathbb{N}$, let $\mathbb{R}^{\mathbb{Z}_{N}}$ denote the set of $N$-tuples of real numbers, together with periodic indexing. That is, for $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$, we may write $\mathbf{X}=(X_{1},X_{2},\ldots,X_{N})$, where the indices represent elements of $\mathbb{Z}_{N}$ with addition modulo $N$. For a vector $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$ and indices $i,j\in\mathbb{Z}_{N}$, we set
@@
where the sums are taken in cyclic order (see Figure 1). We follow the convention that $\mathbf{X}_{(i,i]}=0$.
@@
$\mathcal{P}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=(T(\mathbf{X}_{1},\mathbf{X}_{2}),D(\mathbf{X}_{1},\mathbf{X}_{2})),\quad\text{where}$ (1.2)
$T_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i-1}+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i,j]}}}\Biggr),\quad\text{and}$
$D_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{1,i+1}+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}\Biggr),\quad\text{where}\quad Y_{\ell}=X_{1,\ell+1}-X_{2,\ell}.$
This map is a shifted version of the map studied in [CGS26] (see Appendix A.1). Note that the maps $\mathcal{P},T,D$ depend on the length $N$ of the periodicity; we will subsume dependence on this value in the notation and fix the positive integer $N$, unless otherwise noted.
$\mathcal{P}_{k}(\ldots,\mathbf{X}_{k-1},\mathbf{X}_{k},\mathbf{X}_{k+1},\mathbf{X}_{k+2},\ldots)=(\ldots,\mathbf{X}_{k-1},T(\mathbf{X}_{k},\mathbf{X}_{k+1}),D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2},\ldots).$ (1.3)
Our first main result is the following:
Theorem 1.1.
The operators $\mathcal{P}_{k}$ satisfy the following:
  1. (i)
    For $k\in\mathbb{Z}$, the map $\mathcal{P}_{k}$ is an involution.
  2. (ii)
    For $k\in\mathbb{Z}$,
  3. (iii)
    For |kj|>1|k-j|>1,
    $\mathcal{P}_{j}\mathcal{P}_{k}=\mathcal{P}_{k}\mathcal{P}_{j}.$
The commutativity in Item (iii) of Theorem 1.1 is immediate from the definition. Item (i) was previously proved in [CGS26, Corollary 5.9], although it is stated slightly differently there for a shifted version of the operator $\mathcal{P}$ (see Appendix A.1). In the present paper, we give an alternative, self-contained proof of Item (i) in Section 2.2. Item (ii) is called the braid relation, and it is also proved in Section 2.2. Crucial to the proof is a matrix relation in Proposition 2.7, which is also used to prove Theorem 1.4. This matrix relation comes from adapting the matrix representation of geometric RSK from the work of Noumi and Yamada [NY04] to the periodic case.
The braid relation consists of three equalities ((2.11)-(2.13)), the third of which was previously proved in [CGS26, Proposition 5.13]. From there, we iterate the matrix relation in Proposition 2.7 in two different ways and use an interpretation of the matrix product as a polymer partition function (lemma 3.1) to obtain the other two equalities. Specifically, Lemma 2.8 is the key innovation that allows us to complete the proof.
We call a permutation $\sigma:\mathbb{Z}\to\mathbb{Z}$ a finite permutation if $\sigma$ fixes all but finitely many coordinates. Let $\mathcal{S}_{\mathbb{Z}}$ denote the group of these finite permutations under composition of functions. The relations satisfied by the maps $\mathcal{P}_{k}$ in Theorem 1.1 are the same as those satisfied by the nearest-neighbor transpositions $(k,k+1)$, which generate the group $\mathcal{S}_{\mathbb{Z}}$. Thus, if $\sigma$ is any such finite permutation, if we decompose it in two different ways as
$\sigma=(k_{1},k_{1}+1)(k_{2},k_{2}+1)\cdots(k_{n},k_{n}+1)=(\ell_{1},\ell_{1}+1)(\ell_{2},\ell_{2}+1)\cdots(\ell_{r},\ell_{r}+1),$
then Theorem 1.1 implies that we have the following equality of operators:
$\mathcal{P}_{k_{1}}\mathcal{P}_{k_{2}}\cdots\mathcal{P}_{k_{n}}=\mathcal{P}_{\ell_{1}}\mathcal{P}_{\ell_{2}}\cdots\mathcal{P}_{\ell_{r}}.$
For any finite permutation $\sigma$, we then define the operator $\mathcal{P}_{\sigma}$ by the product above. The operators $\mathcal{P}_{\sigma}$ therefore define a group action of $\mathcal{S}_{\mathbb{Z}}$ on the space $(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$.
@@ 1.2. Polymers on $\mathbb{Z}^{2}$ and periodic environments
$\pi:=\left\{(a_{i},b_{i})_{i\in\llbracket 1,L\rrbracket}\mid(a_{i+1},b_{i+1})-(a_{i},b_{i})\in\{(1,0),(0,1)\}\ \forall 1\leq i<L\right\},$ (1.4)
where $(a_{i},b_{i})\in\mathbb{Z}^{2}\ \forall i$. Here, the notation $\llbracket a,b\rrbracket$ denotes the set of $k\in\mathbb{Z}$ satisfying $a\leq k\leq b$. We also define the set of all up-right paths to be $\Pi_{\mathbb{Z}^{2}}$.
@@
to be the set of paths starting at $(i,j)$ and ending at $(n,m)$, noting that $\Pi_{(i,j),(n,m)}\neq\varnothing$ if and only if $m\geq j$ and $n\geq i$. In this case, L=m+njiL=m+n-j-i for any $\pi$ in $\Pi_{(i,j),(n,m)}$. We say that two paths $\pi_{1},\pi_{2}\in\Pi_{\mathbb{Z}^{2}}$ are non-intersecting if $\pi_{1}\cap\pi_{2}=\varnothing$. We say that a non-intersecting $k$-tuple of paths $(\pi_{1},\ldots,\pi_{k})$ is called a multi-path. See Figure 2 for a visual example.
@@
$\Pi_{U,V}:=\{\pi=(\pi_{1},\dots,\pi_{k})\mid\pi_{i}\in\Pi_{(a_{i},b_{i}),(c_{i},d_{i})}\ \forall i\in\llbracket 1,k\rrbracket,\ \pi_{i}\cap\pi_{j}=\varnothing\ \forall i\neq j\in\llbracket 1,k\rrbracket\}.$ (1.6)
We will often require that there is a unique pairing of initial points in $U$ and end points in $V$ such that a non-intersecting multi-path exists. In other words, we require that if $U$ and $V$ each have $k$ elements, then there does not exist a nontrivial permutation $\varphi$ in the symmetric group on $k$ elements, $\mathcal{S}_{k}$, such that $\Pi_{U,V_{\varphi}}\neq\varnothing$ where $V_{\varphi}:=\bigl((c_{\varphi(i)},d_{\varphi(i)})\bigr)_{i\in\llbracket 1,k\rrbracket}$. We denote the pairs of $k$-tuples of ordered pairs satisfying this property as
$\Psi:=\bigl\{(U,V)\mid U,V\subseteq\mathbb{Z}^{2},\ |U|=|V|,\ \Pi_{U,V}\neq\varnothing,\ \text{ and }\Pi_{U,V_{\varphi}}=\varnothing\ \forall\varphi\in\mathcal{S}_{|U|}\setminus\{\text{Id}\}\bigr\}.$ (1.7)
As an example, if $U=\{(0,0),(2,0)\}$ and $V=\{(2,4),(4,4)\}$ (as depicted in Figure 2), then $(U,V)\in\Psi$. More generally, a pair $(U,V)$ lies in $\Psi$ if and only if $\Pi_{U,V}\neq\varnothing$, the points of $U$ lie along some down-right path, and the points of $V$ lie along a down-right path. See Figure 3 for an example of a pair $(U,V)\notin\Psi$.
@@
[…]
Figure 3. Two multipaths in $\mathbb{Z}^{2}$ from $U:=\{u_{1},u_{2}\}$ (green) to $V:=\{v_{1},v_{2}\}$ (orange), where $(U,V)\notin\Psi$. To see this, note that there exists a multipath from $(u_{1},u_{2})$ to $(v_{1},v_{2})$ (denoted here in blue/thick) as well as a multipath from $(u_{1},u_{2})$ to $(v_{2},v_{1})$ (denoted here in red/thin).
@@
For real-valued weights $\mathbf{X}=\bigl(\mathbf{X}_{i}\in\mathbb{R}^{\mathbb{Z}}\bigr)_{i\in\mathbb{Z}}=\bigl(X_{(i,j)}\in\mathbb{R}\mid i,j\in\mathbb{Z}\bigr)$ corresponding to the vertices of the directed lattice $\mathbb{Z}^{2}$, we define the polymer partition function with inverse temperature $\beta>0$ between the initial point $(i,j)$ and end point $(n,m)$ in $\mathbb{Z}^{2}$ as
$Z^{\beta,\mathbf{X}}(n,m\mid i,j):=\sum_{\pi\in\Pi_{(i,j),(n,m)}}\prod_{(r,\ell)\in\pi}e^{\beta X_{(r,\ell)}}.$ (1.8)
@@
$Z^{\beta,\mathbf{X}}(V\mid U):=\sum_{\pi\in\Pi_{U,V}}\prod_{(r,\ell)\in\pi}e^{\beta X_{(r,\ell)}}.$ (1.9)